3x-1/4 + x/2 = 3/8
1
Expert's answer
2012-10-31T10:03:22-0400
3x-1/4 + x/2 = 3/8
6x-1/2 + x = 3/4
7x=3/4+1/2=(3+2)/4=5/4
x=5/28